what is the length and width if the perimeter is 70 and the width is 11 feet longer than the length?
Answer by waynest(281)

let L = w + 11
P = 2W + 2L
70 = 2W + 2(w + 11)
70 = 2W + 2w + 22
70 - 22 = 4W
48 = 4W
48/4 = W
12 = W
check:
2(12) + 2(12 + 11)= 70
24 + 46 = 70
70 = 70